2.3 The National Safety Council reports the following age breakdown for licensed drivers in the United States Source: Bureau of the Census, Statistical Abstract of the United States 2009, p. 682. Age (Years) Licensed Drivers (Millions) under 20 20-under 25 25-under 35 35-under 45 45-under 55 55-under 65 65 or over 10.72 16.84 40.78 45.59 36.70 23.32 28.76 Identify the following for the 35-under 45 class: (a) frequency, (b) upper and lower limits, (c) width and (d) midpoint.
media%2F7ba%2F7ba3250f-0e4a-4a07-89e3-67
media%2F8f1%2F8f14a007-9c3e-4f79-9dc7-2f
media%2F496%2F49637bd9-8f0c-4ab0-b1fb-15
Answer #1

Problem 2.3

For Class 35 - less than 45

a. Frequency= 45.59 million licensed drivers

b. Upper limit: 45 years

    Lower limit: 35 years

c. Width = Difference in limit = 45 -35 = 10 years

d. Midpoint = upper limit + Lower limit /2 = ( 45+35) /2 = 40 years